OpenStudy (anonymous):

MEDAL GIVEN!!! One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there. @CoolKat16

OpenStudy (anonymous):

@CoolKatt16

OpenStudy (whpalmer4):

Well, you solve two equations in two unknowns: \[2p + 4h = 196\] \[p+h=74\] One equation gives you the number of legs, assuming every person has 2 and every horse has 4. The other gives you the number of heads, assuming every person has 1 and every horse has 1. You'll have to decide if these are good assumptions :-)

OpenStudy (solomonzelman):

lol, he is a weird guy, counting heads and legs instead of .. ?!

OpenStudy (solomonzelman):

anyway. p=person h=horse p+h=5 legs=2 heads 74 heads i'll assume including the horses' heads ok?

OpenStudy (whpalmer4):

Another way to look at it: assume that he only saw horses. If he saw 74 heads, and they are all 4-legged horses, there would have to be 4*74 = 296 legs. That's 100 more legs than he saw. Converting a horse to a human keeps the same number of heads, but reduces the leg count by 4-2=2 legs. 100 extra legs -> 100/2 = 50 humans.